OpenStudy (anonymous):

Kelsey painted one-fourth of her bedroom in three-sevenths of an hour. At this rate, how long would it take her to paint the entire room

jimthompson5910 (jim_thompson5910):

We have this ratio \[\Large \frac{1}{4} \ \text{bedroom} \ : \ \frac{3}{7} \ \text{hour}\] Multiply both parts by 4 to turn the "1/4" into "4/4" which is just 1 \[\Large {\color{red}{4*}}\frac{1}{4} \ \text{bedroom} \ : \ {\color{red}{4*}}\frac{3}{7} \ \text{hour}\] \[\Large \frac{4}{4} \ \text{bedroom} \ : \ \frac{12}{7} \ \text{hour}\] \[\Large 1 \ \text{bedroom} \ : \ \frac{12}{7} \ \text{hour}\]

jimthompson5910 (jim_thompson5910):

12/7 = 1.714 approximately it takes roughly 1.714 hours to paint the whole room

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

Margie calculated that she would spend $175 on school supplies this year. She actually spent $97.50 on school supplies. What is Margie’s percent of error

OpenStudy (anonymous):

77.3% 79.5% 44.2% 4.4%

jimthompson5910 (jim_thompson5910):

have a look at this page https://www.mathsisfun.com/numbers/percentage-error.html

OpenStudy (anonymous):

i get 72.5

jimthompson5910 (jim_thompson5910):

Did you use the formula \[\Large \frac{|\text{Approximate Value} - \text{Actual Value}|}{| \text{Actual Value}|}\times 100\%\] they show it on the page

OpenStudy (anonymous):

oh no i didnt

jimthompson5910 (jim_thompson5910):

They also posted this How to Calculate Here is the way to calculate a percentage error: Step 1: Calculate the error (subtract one value form the other) ignore any minus sign. Step 2: Divide the error by the exact value (we get a decimal number) Step 3: Convert that to a percentage (by multiplying by 100 and adding a "%" sign)

OpenStudy (anonymous):

whats the approximate value

OpenStudy (anonymous):

nvm

jimthompson5910 (jim_thompson5910):

it's the estimate

OpenStudy (anonymous):

i couldnt find the answer can you tell me ill give you a medal

jimthompson5910 (jim_thompson5910):

approximate = 175 (this is the amount she estimates) actual = 97.50 (actual amount she spends)

jimthompson5910 (jim_thompson5910):

does that help?